OpenStudy (anonymous):

-1/2 x + 2 = −x + 7 My answer is 2/5 Am I right?

OpenStudy (anonymous):

Try again.

OpenStudy (anonymous):

5/2?

OpenStudy (anonymous):

Try again

OpenStudy (anonymous):

use order of operations...

OpenStudy (anonymous):

-1/2 x + 2 = −x + 7 -1/2 x + 2 + x = −x + 7 + x 1/2 x + 2 = 7

OpenStudy (anonymous):

I got ya started. Now you finish it

OpenStudy (anonymous):

Is the answer 10

OpenStudy (anonymous):

I didn't do it the same way as him so idk:)

OpenStudy (anonymous):

Yes !!

OpenStudy (anonymous):

How did you dot it @Emah . The only other way would be to add a positive 1/2 x which would mean x would be negative on the right and complicate the problem more than it needs to be. I always try to keep x positive if I can.

OpenStudy (anonymous):

I just put all the same variables on the same side, then simplified it.

OpenStudy (anonymous):

But I got the same answer, I was just trying to figure out the way you did it. @ ChmE